OpenStudy (anonymous):

A circle of radius r has a central angle of θ=15° which intercepts an arc of 27 inches. what is the value of radius r?

OpenStudy (anonymous):

use r=s/θ =27/(pi/12)=27*12/pi =103.09 appx